# MIXRTs: Toward Interpretable Multi-Agent Reinforcement Learning via Mixing Recurrent Soft Decision Trees

### v3

@@ 6.2 Feature Importance
There are several ways of implementing feature importance assignments on SDTs [31]. However, the data point is more susceptible to being perturbed since it is less confident of remaining in the original when there are multiple boundaries for partitioning the space. We utilize decision confidence as a weight for assigning feature importance, which can be positively correlated with the distance from the instance to the decision boundary to relieve the above effects. Therefore, similar to Eq. (8), we weight the confidence probability of reaching the deepest non-leaf level node $j$ as $P^{j}(o_{i}^{t},h_{i}^{t-1})$. Then, the feature importance can be expressed by multiplying the confidence value with each decision node as
@@
Feature importance analysis. After obtaining learned RTCs, we evaluated the feature importance of agents on different maps with Eq. (15). we select three feature properties of the agent to indicate their importance and display the weight of individual credit assignment across different episodes, including the agent’s own health, enemy distance, and relative Y properties. We analyze these properties on three different maps, and the results are shown in Figs. 8(a)8(c), and 8(e), respectively. The horizontal coordinate represents the number of steps in the episode, and two vertical coordinates represent the corresponding value of the agent property and feature importance, respectively. Meanwhile, we visualize the weights $W_{i}$ for each agent $i$ on each scenario, and the weight heatmaps are shown in Figs. 8(b), 8(d) and 8(f). In the attention heatmaps, the steps increase from bottom to top, and the horizontal ordination indicates the agent id.
@@
For 8m, it is essential to achieve a victory that the agents avoid being killed and pay more attention to firepower for killing the enemies. From Fig. 8(a), we observe that allies have similar feature importance on the same attribute. It is worth noting that when an enemy is killed in combat, both the importance of the distance to this enemy and its relative Y will decrease. It suggests that RTCs can capture the skills for accomplishing tasks and generate coping actions when the environment changes. In addition, as shown in Fig. 8(b), we can find that each agent has almost equal attention weights, indicating that each allied Marine plays a similar role on the homogeneous scenario, and MIXRTs almost equally divides $Q_{tot}$ for each agent. MIXRTs performs well with the refined mixing trees since it adjusts the weights of $Q_{i}$ with a minor difference. For the 2s3z scenario, we can find that Stalkers and Zealots each have similar feature importance at different steps, which might be sourced from the fact that they all focus on their health and play an equally important role in the battle. Besides, when the health value of the Zealot $2$ is 00, we also find that it achieves strongly negative importance with blue shades as shown in Fig. 8(d), where the agent does not play an active role in battle since it is killed. It is of interest that the drastic change in the weight of the agents on 2s3z since the replacement of Zealot $2$ by Zealot 00 after its death in battle, coinciding with a rise in the importance of distance and health. Furthermore, we notice that different types of soldiers have different sensitivity to features. For example, Medivac (agent $9$ in Fig. 8(f)) receives more attention during the early stages of the battle, which may be attributed to their different roles involved in combat. In summary, by analyzing the feature importance, we receive meaningful implicit knowledge of the tasks, which can facilitate human understanding of why particular feature importance leads to a certain action.
@@
Perturbing important features. Due to the large number of features involved in the decision-making process, we mask certain features by degradation of performance to measure their importance. For a trained MIXRTs, we first calculate the feature importance at each step through the nodes. Then, we evaluates performance by substituting the most and least important features with zeros, implying that the corresponding nodes in the tree structure are not involved in the calculation. The performance results of win rate are presented in Table V. When perturbing a portion of unimportant features remains relatively stable in performance, which suggests that the critical decision-making process does not heavily rely on the full set of features. Conversely, the perturbing of features deemed important leads to a marked degradation in performance, affirming their significance in the decision-making process. Furthermore, with an increase in the proportion of masked features, the win rate of perturbing critical features precipitously declines to zero, while vice versa does not. This demonstrates the effectiveness of MIXRTs in evaluating the importance of features and allows for the exclusion of a portion of features that are inconsequential to the decision-making process.
@@ 6.3 Stability on Feature Importance
Figure 9: Depicting the stability on fitted MIXRTs through violin plots. Broad sections of the plot denote a higher density distribution, while narrow portions imply a less dense distribution.
@@
The stability of the interpretable model is an important factor reflecting the reliability. To further study the stability, we take a deeper look into MIXRTs itself regarding the assigned feature importance and the action distributions. We conduct analyses with the violin plot to study action distribution among several episodes. With the violin plot, it would be beneficial for us to immediately identify the median feature importance size without having to visually estimate it by integrating the density, which can help us receive strictly more information and analyze the stability of feature importance. We hope this can alleviate the differences in action distributions caused by the different initial states of each episode. As shown in Fig. 9, we select the health property of the agents to indicate the correlation between the underlying action distribution and the assigned feature importance over $32$ episodes. The horizontal coordinate and the vertical coordinate represent the selected action and the importance of the health of the agent itself.
@@
For the 3m scenario, it requires the three allied agents to pay more attention to firepower to kill the enemies with fewer casualties. As shown in Fig. 9, we can observe that the three allies own similar importance among all the actions, which aligns with the common knowledge that homogeneous agents play the same important role during the battles. Besides, we also find that they pay more attention to selecting attack actions, which may mean that it can bring more positive rewards to ensure victory than taking the other actions in this easy homogeneous agent environment. From Figs. 9(b) and 9(c), considerable differences can also be spotted over $32$ episodes on the heterogeneous scenarios 2s3z and MMM2, even though MIXRTs has captured similar skills to win. For the 2s3z scenario, we see Zealots obtaining a higher importance of health than Stalkers, which may require a better winning strategy where Zealots agents serve at the front of combat, killing enemies one after another while protecting the Stalkers to kite the enemy around the map. Furthermore, we notice that similar interesting findings also exist on the MMM2 scenario. For example, Medivac, as a healer unit, receives more attention than other kinds of agents during the battle. This could lead the cooperative team to defeat the enemies more since Medivac uses healing actions to improve the health of its allies. It consistent with the above analysis in Section 6.2.
@@
Overall, the importance of health is slightly higher when the agents are in attacking status than in moving status, and is significantly higher than when they are not operating. Since on most occasions when the agents are attacking, they are also being attacked, the importance of health is more intense. Meanwhile, agents who are killed receive the negative importance of health with a small interquartile range. As far as the stability is concerned, the non-operation of agents also maintains a steady state in the model since most of them own a health value of 00. In summary, even though the different initialized environments change, we could still find the implicit knowledge from MIXRTs by analyzing the feature importance over several episodes.
@@ 6.4 Case Study
Figure 10: Visualization of the evaluation for MIXRTs on 2s3z and MMM2 scenarios in SMAC (time step goes from left to right). For each task, the credit assignment weights for each agent are displayed at the top-right of each frame. The arrow and the colored circle indicate the direction that each moving agent is facing and the central attack range, respectively. For the 2s3z scenario, the agents will receive higher credit when they focus on killing the enemy within its range of attacking, which implies the credits reflecting their contributions to the team. They learn a skill-winning strategy that they collectively wall off the choke point to block enemy attacks from different directions as shown in frames (a) and (b). For the MMM2 scenario, the assigned credit for each unit indicates the important role that Medivac receives the highest credit during the early stage of the battle since it can heal the allied units, and Marauders receive higher credit than Marines for their high damage.
To further show the interpretability of MIXRTs, we present the keyframes on these episodes during the testing stage in Fig. 10. As seen from Fig. 10(a), during the initial stage (step=$10$) of 2s3z, Zealot $2$ and Zealot 00 are participating in the battle, and meanwhile, Zealot $1$ drops teammates. It can be understood that Zealot $2$ and Zealot 00 make contributions to the team and Zealot $1$ does not contribute to the team. Their credit assignment weights $0.23344$, $0.22639$, and $0.16338$ correctly catch the manner of Zealot 00, Zealot $2$ and Zealot $1$, respectively. When Zealot $2$ was killed as shown in Fig. 10(b), its credit assignment weight also quickly dropped to the lowest team level, and meanwhile, its allies’ weight increased. During the battle process, both Stalkers capture high-level skills with stable credit assignment weights when the team is in a combat tactic, where they nearly always attack enemies behind teammates for getting the utmost out of their long-range attack attributes. Besides, Zealot 00 and Zealot $1$ focus on killing the enemy as they obtain higher credit at step $34$. As a result, Zealot 00 and Zealot $1$ contribute more than others and therefore their credits are greater, which implies that the credits reflect contribution to the team. For the MMM2 scenario, we find an interesting strategy to win the battle that the being attacked Medivac will retreat out of being attacked range while the allies move forward to prevent Medivac from being attacked. Its credit assignment weight correctly catches the important role of its team with the highest credit during the early stage of the battle. Besides, Marauders (Agents $7$ and $8$) receive higher credit for their high damage, they get healing blood from the Medivac agent as shown in Figs. 10(f). In a summary, MIXRTs can help us understand the behaviors of agents and their contribution to the team via linear credit assignments on different complicated scenarios.
@@ 6.5 User Study
Our final evaluation investigates the effect of the interpretability provided by MIXRTs for non-experts. Following Silva et al. [28, 55], we have similarly conducted a user study, showing participants trained policies on SMAC scenarios. They are tasked with identifying whether these agents’ credit assignments are reasonable given a set of observation inputs and keyframes [^2] Due to the overwhelmingly large features, we only interpret credits to alleviate frustration for non-expert participants (similar to Fig.  10). We compare the credit assignment rationality of MIXRTs, VDN, and QMIX. We then present our results to $20$ participants, instructing each of them to assign a $1\sim 5$ helpfulness score with a set of Likert scales to every result. The Likert scale guideline is provided on each score for justified scoring as follow
  • Score $1$ for misleading credits that impair the decision-making process.
  • Score $2$ for uninformative credits that neither aid nor hinder human understanding.
  • Score $3$ for deviated but somewhat related credits that provide slight assistance.
  • Score $4$ for strongly related keyframes but slightly inaccurate, helpful.
  • Score $5$ for well-aligned credits that contain all information in combat, exceptionally helpful.
@@
Figure 11: Results from our user study. Higher Likert ratings are better, and lower time taken is better.
Results of our user study are shown in Fig. 11. Consistent with intuition and previous metrics, human evaluators reveal a positive estimation of $4.0485$ helpfulness rating. The rating indicates our MIXRTs generally offers a moderate enhancement to interpretability. Notably, this result substantially outperforms QMIX’s $2.9979$ helpfulness and VDN’s $3.2640$ helpfulness, both of which means vague and deviate related content. We conduct an ANOVA, which yields that $F(2,57)=33.40,p<0.0001$ for the Likert scale ratings. It indicates that different mixers significantly effect the interpretability and usability of credit assignment. A Tukey’s HSD post-hoc test shows that MIXRTs vs. QMIX is $t=10.26,p<0.0001$, and MIXRTs vs. VDN is $t=7.67,p<0.0001$. We also test participants’ time in Fig. 11 and show that MIXRTs takes the least amount of time compared to the other methods. An interesting observation from the study participants is that they would have abandoned the task if what they knew was the credit generated by QMIX. These results support the hypothesis that our model significantly outperforms other approaches in reducing human frustration and improving interpretability. These empirical insights lend robust support to the supposition that our proposed model markedly excels in diminishing human frustration and enhancing interpretability.
@@ 7 Conclusion
This paper presented MIXRTs, a novel interpretable MARL mixing architecture based on the SDTs with recurrent structure. MIXRTs allows end-to-end training in a centralized fashion and learns to linearly factorize a joint action-value function for executing decentralized policies. The empirical results show that MIXRTs not only enjoy good model interpretability and scalability while maintaining competitive learning performance. Our approach captures the implicit knowledge of the challenging tasks with a transparent model and facilitates understanding of the learned domain knowledge and how input states influence decisions, which paves the application process in high-stake domains. Our linear mixing component demonstrates the possibility to produce the credit assignment to analyze what role each agent plays in its allies. Our work motivates future work toward building more interpretable and explainable multi-agent systems. In future work, we will try to analyze the evaluation process, and approach a more trustworthy interpretation of the feature importance estimation, aiming at further minimizing the cost of human effort for understanding the decision process.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 62006111 and Grant 62073160.

### v4

@@ 6.2 Feature Importance
There are several ways of implementing feature importance assignments on SDTs [32]. However, the data point is more susceptible to being perturbed since it is less confident of remaining in the original when there are multiple boundaries for partitioning the space. To mitigate this issue, we utilize decision confidence as a weighting factor in feature importance assignment, which can be positively correlated with the distance from the instance to the decision boundary to relieve the above effects. Therefore, similar to Eq. (5), we weight the confidence probability of reaching the deepest non-leaf level node $j$ as $P^{j}(o_{i}^{t},h_{i}^{t-1})$. The feature importance can be expressed by combining these confidence values with the weights of each decision node
@@
Feature importance analysis. After obtaining learned RTCs, we evaluated the feature importance of agents on different maps with Eq. (13). We select three agent feature properties to indicate their importance and display the weights of individual credit assignments across different episodes, including the agent’s health, enemy distance, and relative Y properties. We analyze these properties on three different maps, and the results are shown in Figs. 8(a)8(c), and 8(e), respectively. The horizontal coordinate represents the number of steps in the episode, and the two vertical coordinates represent the corresponding value of the agent property and feature importance, respectively. Meanwhile, we visualize the weights $W_{i}$ for each agent $i$ on each scenario, and the weight heatmaps are shown in Figs. 8(b), 8(d) and 8(f). In the attention heatmaps, the steps increase from bottom to top, and the horizontal ordination indicates the agent id.
@@
For the 8m scenario, it is essential to achieve a victory that the agents avoid being killed and pay more attention to firepower for killing the enemies. From Fig. 8(a), we observe that allies agents have similar feature importance on the same attributes. It is worth noting that when an enemy is killed in combat, both the importance of the distance to this enemy and its relative Y will decrease. This indicates that RTCs can capture the skills for accomplishing tasks and generating coping actions in response to environmental changes. In addition, as shown in Fig. 8(b), we can find that each agent has almost equal attention weights, indicating that each allied Marine plays a similar role in the homogeneous scenario, and MIXRTs almost equally divides $Q_{tot}$ for each agent. MIXRTs perform well with the refined mixing trees since they adjust the weights of $Q_{i}$ with a minor difference. For the 2s3z scenario, we observe that Stalkers and Zealots have similar feature importance at different steps, which might be sourced from the fact that they all focus on their health and play an equally important role in the battle. Furthermore, when the health value of the Zealot $2$ drops to 00, we find that it exhibits strongly negative importance with blue shades as shown in Fig. 8(d), where the agent does not play an active role in battle since it is killed. Interestingly, on the 2s3z scenario, there is a significant shift in the weight assignments among the agents. Following the elimination of Zealot $2$ during the battle and its subsequent replacement by Zealot 00, we observe a marked increase in the importance of the distance and health features. Moreover, we observe that different types of soldiers have different sensitivity to features. For example, Medivac (agent $9$ in Fig. 8(f)) receives more attention during the early stages of the battle, which may be attributed to its unique role as a support unit in combat. In summary, by analyzing the feature’s importance, we receive meaningful implicit knowledge about the tasks, which can facilitate human understanding of why a particular feature’s importance leads to a specific action.
@@
Perturbing important features. Due to the large number of features involved in the decision-making process, we mask certain features by degradation of performance to measure their importance. For trained MIXRTs, we first calculate the feature importance at each step through the nodes. Then, we evaluate performance by substituting the most and least important features with zeros, implying that the corresponding nodes in the tree structure are not involved in the calculation. The performance results of the win rate are presented in Table V. When perturbing a portion of unimportant features, the performance remains relatively stable, which suggests that the critical decision-making process does not heavily rely on the full set of features. On the contrary, perturbing the features consider red important leads to a significant decrease in performance, affirming the importance of these features in the decision-making process. Furthermore, as the proportion of masked features increases, the win rate declines precipitously to zero when critical features are perturbed, whereas it does not when unimportant features are masked. This demonstrates the effectiveness of MIXRTs in evaluating the importance of features and allows for excluding a portion of features that are inconsequential to the decision-making process.
@@ 6.3 Stability on Feature Importance
Figure 9: Depicting the stability of fitted MIXRTs through violin plots. Broad sections of the plot denote a higher density distribution, while narrow portions imply a less dense distribution.
@@
The stability of an interpretable model is an important factor reflecting reliability. To further investigate stability, we delved into the feature importance assignments and action distributions of the MIXRTs. We conduct analyses with the violin plot to study action distribution over several episodes. The violin plot helps us to immediately identify the median feature importance without the need to visually estimate it by integrating the density, thereby providing more precise information to analyze the stability of feature importance. We hope this can alleviate the differences in action distributions caused by the different initial states of each episode. As shown in Fig. 9, we select the health property of the agents to indicate the correlation between the underlying action distribution and the assigned feature importance over $32$ episodes. The horizontal coordinate and vertical coordinate represent the selected action and the importance of agent health, respectively.
@@
On the 3m scenario, three allied agents focus on firepower to kill the enemies with fewer casualties. As shown in Fig. 9, the agents exhibit similar importance across all actions, which aligns with the common knowledge that homogeneous agents play the same important role during the battles. Besides, we find that they prioritize selecting attack actions, as these yield more positive rewards to ensure victory in this easy homogeneous environment. From Figs. 9(b) and 9(c), considerable differences can also be spotted over $32$ episodes on the heterogeneous scenarios 2s3z and MMM2, even though MIXRTs have captured similar skills to win. For the 2s3z scenario, Zealots display a higher importance of health than Stalkers, which may require a better winning strategy where Zealots agents serve at the front of combat, killing enemies one after another while protecting the Stalkers to kite the enemy around the map. Furthermore, we notice that similar interesting findings also exist in the MMM2 scenario. For example, Medivac, serving as a healer unit, receives more attention than other kinds of agents during the battle. This focus on Medivac likely enhances the team’s effectiveness, as it uses healing actions to improve the health of its allies, which is consistent with the analysis in Section 6.2.
@@
Overall, the importance of health is slightly higher when the agents are in attacking status than in moving status, and it is significantly higher than when not operating. Since agents are often being attacked while they are attacking, the importance of health is more intense. Meanwhile, agents who are killed receive the negative importance of health with a small interquartile range. Regarding stability, agents that are not operating maintain a steady state in the model, as their health value is 00. In summary, even when initial environmental conditions vary, we can still find implicit knowledge from MIXRTs by analyzing the feature importance over several episodes.
@@ 6.4 Case Study
Figure 10: Visualization of the evaluation for MIXRTs on 2s3z and MMM2 scenarios in SMAC (time step goes from left to right). For each task, the credit assignment weights for each agent are displayed at the top-right of each frame. The arrow and the colored circle indicate the direction that each moving agent is facing and the central attack range, respectively. For the 2s3z scenario, the agents will receive higher credit when they focus on killing the enemy within its range of attacking, which implies the credits reflect their contributions to the team. As shown in figures (a) and (b), they learn a skill-winning strategy of blocking enemy attacks launched from different directions. For the MMM2 scenario, the assigned credit for each unit indicates the important role that Medivac receives the highest credit during the early stage of the battle since it can heal the allied units, and Marauders receive higher credit than Marines for their high damage.
To further demonstrate the interpretability of MIXRTs, we present the keyframes on these episodes during the testing stage in Fig. 10. As observed from Fig. 10(a), during the initial stage (step=$10$) of the 2s3z scenario, Zealot $2$ and Zealot 00 are actively participating in the battle, while Zealot $1$ drops teammates. It can be understood that Zealot $2$ and Zealot 00 make contributions to the team, whereas Zealot $1$ is not significantly involved. Their credit assignment weights $0.23344$, $0.22639$, and $0.16338$ correctly catch the manner of Zealot 00, Zealot $2$ and Zealot $1$, respectively. When Zealot $2$ was killed as shown in Fig. 10(b), its credit assignment weight also quickly dropped to the lowest team level, while the weights of its allies increased. During the battle process, both Stalkers capture high-level skills with stable credit assignment weights when the team is in a combat tactic, where they nearly always attack enemies behind teammates to get the utmost out of their long-range attack attributes. Besides, Zealot 00 and Zealot $1$ focus on killing the enemy, obtaining higher credit assignment weights at step $34$. As a result, Zealot 00 and Zealot $1$ contribute more than the others with greater credits, indicating that the credit assignments accurately reflect their contributions to the team. For the MMM2 scenario, we find an interesting strategy to win the battle: when the Medivac is under attack, it retreats out of the enemy’s range while its allies advance to shield it from further attacks. Its credit assignment weight correctly catches the important role of its team with the highest credit during the early stage of the battle. Besides, Marauders (Agents $7$ and $8$) receive higher credits due to their significant damage output, they receive healing blood from the Medivac agent as shown in Figs. 10(f). In summary, MIXRTs can help us understand the behaviors of agents and their contributions to the team through linear credit assignments on different complex scenarios.
@@ 6.5 User Study
Our final evaluation investigates the effect of the interpretability provided by MIXRTs for non-expert users. Following the methodology of Silva et al. [29, 55], we design a user study where participants are shown policies trained on SMAC scenarios. They are tasked with identifying whether these agents’ credit assignments are reasonable given a set of observation inputs and keyframes [^2] Due to the overwhelmingly large features, we only interpret credits to alleviate frustration for non-expert participants (similar to Fig.  10). We compare the credit assignment rationality among MIXRTs, VDN, and QMIX. We present our results to $20$ participants, instructing each to assign a helpfulness score from $1$ to $5$ using a Likert scale to each result. The Likert scale guideline is provided on each score for justified scoring as
  • Score $1$: misleading credits that impair the decision-making process.
  • Score $2$: uninformative credits that neither aid nor hinder human understanding.
  • Score $3$: deviated but somewhat related credits that provide slight assistance.
  • Score $4$: strongly related keyframes but slightly inaccurate, helpful.
  • Score $5$: well-aligned credits that contain all information in combat, exceptionally helpful.
@@
Figure 11: Results of our user study, where a higher Likert rating or lower time taken is better.
The results of our user study are shown in Fig. 11. Consistent with intuition and previous metrics, human evaluators reveal our MIXRTs a positive helpfulness rating of $4.05$. The rating indicates that our MIXRTs generally offer a moderate enhancement to interpretability. Notably, this result substantially outperforms QMIX’s helpfulness rating of $2.9979$ and VDN’s rating of $3.2640$, both of which suggest vague and deviate-related content. We conduct an ANOVA on the Likert scale ratings, yielding $F(2,57)=33.40,p<0.0001$. This indicates that different mixing methods significantly affect the interpretability and usability of credit assignments. A Tukey’s HSD post-hoc test showed significant differences between MIXRTs and QMIX ($t=10.26$, $p<0.0001$) and between MIXRTs and VDN ($t=7.67$, $p<0.0001$). We also measure the time participants took to evaluate each method, as shown in Fig. 11. MIXRTs required the least amount of time compared to the other methods. An interesting observation from the study participants is that some participants expressed they would have abandoned the task if they had known the credits were generated by QMIX. These results support the hypothesis that our model significantly outperforms other approaches in reducing human frustration and improving interpretability. These empirical insights strongly support the supposition that our proposed model markedly excels in diminishing human frustration and enhancing interpretability.
@@ 7 Conclusion
This paper presented MIXRTs, a novel interpretable MARL mixing architecture based on the SDTs with recurrent structure. MIXRTs allow end-to-end training in a centralized fashion and learn to linearly factorize a joint action-value function for executing decentralized policies. The empirical results show that MIXRTs provide not only good model interpretability but also competitive learning performance. Our approach captures the implicit knowledge of the challenging tasks with a transparent model and facilitates understanding of the learned domain knowledge and how input states influence decisions, which paves the application process in high-stake domains. Besides, the linear mixing component demonstrates the possibility of generating credit assignments to analyze what role each agent plays among its allies. Our work motivates future work toward building more interpretable and explainable multi-agent systems. In future work, we will attempt to analyze higher-level strategies and provide more reliable interpretation estimates, aiming to further reduce the human effort required to understand the decision-making process.